OpenStudy (anonymous):

How do you go from x = a(x + 10) to x = 10a /(1-a) ?

OpenStudy (anonymous):

x = ax + 10a x - ax = 10a x(1-a) =10a -> x = 10a(1-a)

OpenStudy (anonymous):

of course! factoring w/ a 1! thank u